OpenStudy (anonymous):

find the derivative of the function y=ln(2x^8+9x)^9/7

OpenStudy (jdoe0001):

use the chain rule \(\bf \large { \cfrac{d}{dx}\left[{\color{brown}{ ln({\color{blue}{ 2x^8+9x}})^{\frac{9}{7}}}}\right]\implies \cfrac{d}{dx}[{\color{brown}{ \square^{\frac{9}{7}} }}]\cdot \cfrac{d}{dx}[{\color{blue}{ \square }}] }\)